# Keeping Medical AI Healthy: A Review of Detection and Correction Methods for System Degradation

### v1

@@ 5.1.1 Distance-Based Methods
Maximum Mean Discrepancy (MMD). MMD is a non-parametric distance metric used to compare two probability distributions based on their samples. It measures the difference between the mean embeddings of the source and target distributions in a reproducing kernel Hilbert space (RKHS). A larger MMD value indicates a greater discrepancy between the two distributions. MMD is especially well-suited for high-dimensional data. In the study [45], the authors utilize a pre-trained autoencoder to extract latent features from chest X-ray images in both source and target datasets. Then they apply MMD to compare the latent feature distributions between the datasets. A significant MMD score signals data drift, effectively identifying subtle shifts such as those introduced by COVID-19. In [46], a Gaussian kernel-based MMD is used for detecting covariant shift of medical datasets and shows good performance.
@@
The study [47] proposes a statistical method to detect domain shifts in medical imaging (e.g., pathology) for deep learning models. Features are first extracted from pathology images using a CNN, and shifts are quantified by comparing feature distributions in the latent space using metrics like Wasserstein Distance. Validated on datasets like CAMELYON17, the method shows a strong link between representation shifts and model performance degradation.
Kullback-Leibler (KL) Divergence. KL Divergence is a fundamental concept from information theory that measures how one probability distribution differs from another reference distribution. Specifically, it quantifies the expected extra information needed to represent data from the target distribution using a model trained on the source distribution. In simple terms, it reflects how inefficient it is to assume the data follows one distribution when it actually follows another. The study [47] also uses KL Divergence to detect domain shifts in medical imaging datasets (e.g., pathology).
Jensen-Shannon (JS) Divergence. It is a symmetric and smoothed variant of KL divergence. It measures the similarity between two probability distributions. In data shift detection, JS Divergence is commonly used to compare feature or prediction distributions between a source (training) and target (test) dataset. In a study analyzing temporal shifts in healthcare data [48], JS divergence is used to compare data shift over time in two datasets: MIMIC-IV and a Brazilian COVID-19 registry. The result shows its robustness for monitoring temporal shifts in real-world healthcare data.
@@ 5.1.2 Hypothesis Test Methods
These methods assess whether two samples come from the same distribution using formal statistical tests. They are typically applied to low-dimensional or 1D data where raw distributional differences are directly measurable. These methods output p-values, enabling principled, threshold-free decisions. Their main strengths are simplicity, interpretability, and strong theoretical grounding, making them effective when assumptions are met.
Kolmogorov-Smirnov (KS) Test. KS test is a non-parametric method for comparing two 1D continuous distributions by evaluating the maximum difference between their cumulative distribution functions (CDFs). It outputs a test statistic and a p-value to assess whether the distributions differ significantly. The CheXstray framework [49] applies KS test to monitor real-time data drift in healthcare by comparing DICOM metadata, image features, and model outputs to a reference dataset, flagging shifts that may affect performance. McUDI [50] is an unsupervised method that selects key features uses the KS test to detect distribution changes between training and test sets. Similarly, [51] applies KS test-based drift detection and enhances it with periodic expert label queries after drift is detected, enabling more accurate performance monitoring.
Chi-Square Test. This is a basic statistical method commonly used for categorical data, such as gender, hospital unit, or diagnosis code. The idea is to compare how often each category appears in two datasets. If the difference in frequencies is too large, the test signals a potential shift. In [49], chi-square test is applied to DICOM metadata, such as gender and projection, to detect categorical distribution shifts in real-time imaging data monitoring. Chi-square test is simple, fast, and interpretable, making it useful in clinical settings. But it is limited to categorical features and cannot handle continuous variables directly.
@@ 5.1.3 Sequential Statistical Methods
These methods aim to detect distributional shifts over time in streaming or temporally ordered data. Unlike batch comparison methods, they analyze data in a continuous fashion, often by monitoring summary statistics or deviations from a baseline. These methods are well-suited for detecting gradual drift in real-time settings, where data arrives sequentially.
Cumulative Sum (CUSUM). This algorithm [52] is a statistical monitoring method for detecting shifts over time by tracking small, incremental deviations from an expected value. When the cumulative sum exceeds a set threshold, it flags a potential shift. This approach is effective for catching subtle, gradual changes and is commonly used in quality control and performance monitoring. The study [53] investigates AI performance degradation on two audio COVID-19 datasets using the CUSUM algorithm to monitor changes in the MMD statistic over time. By tracking the cumulative sum of MMD values across sliding data windows, it detects sustained deviations that signal drift.
Statistical Process Control (SPC). SPC is a versatile and broadly applicable framework for monitoring healthcare data over time. Originally developed for manufacturing, SPC is now widely used in healthcare and machine learning applications to detect shifts or anomalies in data distributions or model performance. It encompasses a range of statistical tools, including control charts, Exponentially Weighted Moving Averages (EWMA), and Cumulative Sum (CUSUM). The study [54] uses SPC to monitor data shifts and detect out-of-distribution (OOD) instances in radiology datasets. It combines control charts with CUSUM to track changes in feature distribution in real time. This method effectively identifies shifts across imaging modalities (e.g., chest X-rays vs. other radiographs) and demographic groups, highlighting SPC’s robustness for monitoring data and model performance in dynamic clinical settings.
@@ 5.2.1 Supervised Methods
Supervised methods detect data shift by training a domain classifier (e.g., logistic regression or support vector machine) to distinguish between source and target samples [46]. The classifier’s performance, measured by AUC, reflects distributional difference: AUC near 0.5 suggests no shift, while a high AUC indicates a clear separation. This approach, also known as Classifier Two-Sample Test (C2ST) [55, 56, 57], is simple, flexible, and effective in high-dimensional data.
The study [58] applies this method to monitor an AI system developed for diabetic retinopathy. Their classifier detected shifts related to image quality, co-morbidities, and demographics, outperforming traditional statistical tests but requiring larger training datasets. [59] uses an adversarial learning approach, training a discriminator to detect shifts between datasets. Given that real-world data shifts can emerge gradually or sporadically, domain classifiers should ideally be updated in an online fashion to maintain reliable shift detection over time [60].
@@ 5.2.2 Unsupervised Methods
These methods do not require labeled data, and often use autoencoders trained to reconstruct inputs from a source distribution [61]. When applied to target data, a significant increase in reconstruction error (e.g., mean squared error (MSE)) can indicate distributional shift. In addition to reconstruction error, these methods can monitor changes in the distribution of latent feature distributions to detect more subtle or structural shifts. For example, [62] used convolutional autoencoders to detect drift in breast cancer mammography data by analyzing both reconstruction error and latent space changes.
@@ 5.3 Feature-Based Methods
These methods focus on detecting the changes in the relationship between input features and model predictions over time. Instead of monitoring raw data or outputs alone, they analyze shifts in feature contributions to the model’s decision-making process, capturing changes in either the input distribution ($p(x)$) or the conditional distribution ($p(y|x)$). This is typically achieved using explainability tools like SHAP [63, 64] or by training auxiliary classifiers to identify which features are driving the observed shift [65]. The study [66] analyzes a model trained on pre-COVID emergency department data and observes performance degradation during COVID. By tracking SHAP values over time, they identified shifts in the importance of respiratory-related features, revealing clinically meaningful drift. In the “DataFix” framework [59], a component called DF-Locate uses a binary classifier to distinguish between source and target data, using feature importance scores to identify the variables most responsible for the shift.
These methods are especially useful for root cause analysis, as they offer both detection and interpretability, which is critical in complex systems like healthcare [67].
5.4 Anomaly Detection
While data shift detection focuses on tracking gradual or long-term changes in data distribution, anomaly detection (AD) targets short-term, instance-level abnormalities, i.e., rare, unexpected, or extreme inputs that deviate significantly from the training distribution [68, 69]. Also known as outlier or out-of-distribution (OOD) detection [70, 71, 72], AD is crucial for identifying high-risk instances that may not be captured by shift detection alone.
AD methods operate by flagging inputs that fall outside the model’s learned data manifold, and can be used to trigger rejection, uncertainty estimation, or real-time alerts. This function is especially important in safety-sensitive settings such as clinical decision support and real-time healthcare monitoring [73, 74, 75, 76], where even a single abnormal input may lead to harmful outcomes. AD plays two key roles: (1) as a pre-filter to exclude extreme outliers before shift detection, and (2) as a standalone safeguard for identifying early or localized system failures. For instance, [77] shows that anomaly detection improves the reliability and safety of healthcare AI systems by filtering out unreliable inputs and preventing erroneous predictions.
@@ 6.1 Performance-Based Methods
These methods detect model drift by tracking changes in metrics such as accuracy, error rate, F1 score, or ROC-AUC using labeled target data [45]. They typically compare performance across time windows or batches to identify significant degradation.
A classic example is [78], which monitors online error rates and triggers drift alarms when errors exceed a threshold, prompting model retraining or recalibration. This approach is model-agnostic and supports various algorithms. [79] applied this method to clinical prediction models, using the Wilcoxon test to detect degradation in terms of ROC-AUC over time. They showed that even high-performing models can experience silent drift in dynamic healthcare environments.
@@ 6.2 Model Output-Based Methods
These methods detect drift by analyzing changes in predicted outputs, such as softmax probabilities or confidence scores without relying on ground-truth labels [45, 80, 81]. These outputs serve as compact proxies for the model’s internal decision process, and shifts in their distribution can indicate performance degradation. Such methods provide a scalable, label-free, and model-agnostic approach for real-time monitoring. They are especially valuable in healthcare, where labeled data are often delayed or unavailable.
@@ 6.2.1 Distance-Based Output Monitoring
These methods are especially effective in multi-class settings, capturing subtle, class-specific changes by leveraging the full output vector, providing more detailed insights into models’ behavior.
@@ 6.2.2 Score-Based Output Monitoring
The study [83] uses maximum softmax probability to flag prediction errors and distribution shifts. Evaluated on both vision and natural language processing (NLP) datasets, the method shows that misclassified or OOD inputs tend to have lower softmax scores, making it a widely adopted baseline for OOD and drift detection. Energy-based methods [84] extend this idea by computing a log-sum-exp energy score from logits, which captures overall activation levels and is sensitive to subtle uncertainties.
@@ 6.3 Auxiliary Model-Based Method
These methods are modular and flexible, preserving the integrity of the primary model while enabling auxiliary modules to be trained independently or adapted to dynamic deployment environments.
@@ 6.4 Calibration Drift Detection
Calibration drift occurs when a model’s predicted probabilities no longer align with actual outcome frequencies, even if overall accuracy remains stable. For example, if a model predicts a class with 0.8 confidence, that prediction should be correct about 80% of the time. In dynamic clinical settings, this alignment can degrade over time.
@@ 7.2 Shift Type Disentanglement Method
Accurate identification of the specific type of distribution shift is essentile for effective RCA. The study [90] proposes a framework that differentiates between label shift, covariate shift, and hybrid shift. The approach uses model output distributions to detect label shifts and leverages feature-based analysis via self-supervised encoders to detect covariate shifts. The combined “Duo detector” which integrates both techniques, achieves superior performance across various shift types. Evaluated on large-scale datasets from chest X-rays, mammography, and retinal fundus images, the method demonstrates strong generalizability.
The key takeaway is that each shift type necessitates talored detection and correction strategies. For example, label shift often requires output reweighting, whereas covariate shift is better addressed through feature-level adjustments.
@@ 8.1.1 Unsupervised Domain Adaptation (UDA)
It assumes access to labeled data in the source domain and unlabeled data in the target domain. The goal is to align the latent representations across domains without requiring labels from the target side.
In the study [53], UDA is implemented to correct for data drift in COVID-19 detection models using audio data. After drift detection, UDA minimizes the MMD between the source (development) and target (post-development) data distributions. This process aligns the feature representations of the two domains without requiring labeled data from the target domain. Experimental result validates UDA’s potential to maintain robust performance in evolving healthcare scenarios.

### v2

@@ 5.1.1 Distance-Based Methods
Maximum Mean Discrepancy (MMD). MMD is a non-parametric distance metric used to compare two probability distributions based on their samples. It measures the difference between the mean embeddings of the source and target distributions in a reproducing kernel Hilbert space (RKHS). A larger MMD value indicates a greater discrepancy between the two distributions. MMD is especially well-suited for high-dimensional data. In the study [52], the authors utilize a pre-trained autoencoder to extract latent features from chest X-ray images in both source and target datasets. Then they apply MMD to compare the latent feature distributions between the datasets. A significant MMD score signals data shift such as those introduced by COVID-19. In [53], a Gaussian kernel-based MMD is used for detecting covariant shift of medical datasets and shows good performance.
@@
The study [51] proposes a statistical method to detect domain shifts in medical imaging (e.g., pathology) for deep learning models. Features are first extracted from pathology images using a CNN, and shifts are quantified by comparing feature distributions using Wasserstein Distance. Validated on datasets like CAMELYON17, the method shows a strong link between data shift and model performance degradation.
Kullback-Leibler (KL) Divergence. KL Divergence is a fundamental concept from information theory that measures how one probability distribution differs from another reference distribution. Specifically, it quantifies the expected extra information needed to represent data from the target distribution using a model trained on the source distribution. In simple terms, it reflects how inefficient it is to assume the data follows one distribution when it actually follows another. The study [51] uses KL Divergence to detect data shift in medical imaging datasets (e.g., histopathology).
Jensen-Shannon (JS) Divergence. It is a symmetric and smoothed variant of KL divergence. It measures the similarity between two probability distributions. In data shift detection, JS Divergence is commonly used to compare feature or prediction distributions between a source (training) and target (test) dataset. In a study analyzing temporal shifts in healthcare data [54], JS divergence is used to compare data shift over time in two datasets: MIMIC-IV and a Brazilian COVID-19 registry. The result shows its robustness for monitoring temporal drift in real-world healthcare data.
@@ 5.1.2 Hypothesis Test Methods
These methods assess whether two samples come from the same distribution using formal statistical tests. They are typically applied to low-dimensional or 1-D data where raw distributional differences are directly measurable. These methods output p-values, enabling principled, threshold-free decisions. Their main strengths are simplicity, interpretability, and strong theoretical grounding, making them effective when assumptions are met.
Kolmogorov-Smirnov (KS) Test. KS test is a non-parametric method for comparing two 1-D continuous distributions by evaluating the maximum difference between their cumulative distribution functions. It outputs a test statistic and a p-value to assess whether the distributions differ significantly. The CheXstray framework [55] applies KS test to monitor real-time data shift in healthcare by comparing DICOM metadata, image features, and model outputs to a reference dataset. McUDI [56] is an unsupervised method that selects key features uses the KS test to detect distribution changes between training and test sets. Similarly, [57] applies KS test to drift detection and enhances it with periodic expert label queries after drift is detected, enabling more accurate performance monitoring.
Chi-Square Test. This method is commonly used for categorical data, such as gender, hospital unit, or diagnosis code. The idea is to compare how often each category appears in two datasets. If the difference in frequencies is too large, the test signals a potential shift. In [55], chi-square test is applied to DICOM metadata (e.g., gender) to detect categorical distribution shifts in imaging data monitoring. Chi-square test is simple, fast, and interpretable, but limited to categorical features and cannot directly handle continuous variables.
@@ 5.1.3 Sequential Statistical Methods
These methods detect distributional shifts in streaming or temporally ordered data by continuously monitoring summary statistics or deviations from a baseline, making them well-suited for gradual drift detection in real-time settings.
Cumulative Sum (CUSUM). CUSUM [58] is a statistical monitoring method for detecting data drift by tracking small, incremental deviations from an expected value. When the cumulative sum exceeds a set threshold, it flags a potential shift. This approach is effective for catching subtle, gradual changes and is commonly used in quality control and performance monitoring. The study [59] investigates AI performance degradation on two audio COVID-19 datasets using the CUSUM algorithm to monitor changes in the MMD statistic over time. By tracking the cumulative sum of MMD values across sliding data windows, it detects sustained deviations that signal shift.
Statistical Process Control (SPC). SPC is a versatile and broadly applicable framework for monitoring healthcare data over time. Originally developed for manufacturing, SPC is now widely used in machine learning applications to detect shifts/anomalies in data distributions or model performance. It encompasses a range of statistical tools, including control charts, exponentially weighted moving averages (EWMA), and cumulative sum. Zamzmi et al.[60] applied SPC with CUSUM to monitor feature distribution shifts in radiology data, effectively detecting out-of-distribution cases across imaging modalities and demographic groups, demonstrating SPC’s robustness in dynamic clinical settings.
@@ 5.2.1 Supervised Methods
These methods detect data shift by training a domain classifier (e.g., logistic regression) to distinguish between source and target samples [53]. The classifier’s performance, measured by AUC, reflects distributional difference: AUC near 0.5 suggests no shift, while a high AUC indicates a clear separation. This approach, also known as Classifier Two-Sample Test (C2ST) [61, 62, 63], is simple, flexible, and effective in high-dimensional data.
The study [64] applies this method to monitor an AI system developed for diabetic retinopathy. Their classifier detects shift related to image quality, co-morbidities, and demographics, outperforming traditional statistical tests but requiring larger training datasets. [65] uses an adversarial learning approach, training a discriminator to detect shifts between datasets. Given that real-world data shift can emerge gradually, domain classifiers should ideally be updated in an online fashion to maintain reliable shift detection over time [66].
@@ 5.2.2 Unsupervised Methods
These methods do not require labeled data, and often use autoencoders trained to reconstruct inputs from a source distribution [67]. When applied to target data, a significant increase in reconstruction error (e.g., mean squared error) indicates data shift. In addition to reconstruction error, these methods can monitor changes in the distribution of latent feature distributions to detect more subtle or structural shifts. For example, [68] uses convolutional autoencoders to detect drift in breast cancer mammography data by analyzing both reconstruction error and latent space changes.
@@ 5.3 Feature-Based Methods
These methods focus on detecting the changes in the relationship between input features and model predictions over time. Instead of monitoring raw data or outputs alone, they analyze shifts in feature contributions to the model’s decision-making process, capturing changes in either the input distribution ($p(x)$) or the conditional distribution ($p(y|x)$). This is typically achieved using explainability tools like SHAP [69, 70] or by training auxiliary classifiers to identify which features are driving the observed shift [71]. The study [72] analyzes a model trained on pre-COVID emergency department data and observes performance degradation during COVID. By tracking SHAP values over time, they identify shifts in the importance of respiratory-related features, revealing clinically meaningful drift. The study [65] trains a binary classifier to distinguish source from target data and uses feature importance scores to identify variables driving the shift.
These methods are also very useful for root cause analysis, as they offer both detection and interpretability, which is critical in complex systems like healthcare [73].
5.4 Out-of-Distribution Detection
While data shift detection tracks distribution changes, out-of-distribution (OOD) detection focuses on instance-level abnormalities, i.e., rare, unexpected, or extreme inputs that deviate from the training distribution [74]. OOD detection, also referred to as anomaly or outlier detection, plays a vital role in identifying high-risk cases that may trigger model failures before widespread drift occurs.
Traditional OOD detection methods use model confidence signals like Maximum Softmax Probability (MSP) [75], but often fail on high-dimensional, heterogeneous clinical data with poorly calibrated confidence scores. Recent advances have introduced more robust approaches: 1) energy- and density-based methods [76] that better capture data geometry and uncertainty; 2) contrastive and representation learning [77] that improve in–out separability in embedding space; and 3) diffusion-based methods [78] that yield richer probabilistic estimates for complex or multimodal healthcare data.
Domain-specific methods have also emerged, such as NERO [79], which enables explainable OOD detection for gastrointestinal imaging, and MultiOOD [80], which extends OOD evaluation across text, image, and tabular modalities. However, OOD benchmarking in healthcare remains limited, as general frameworks like OpenOOD [81] rarely capture real-world clinical variability.
OOD detection plays two key roles: 1) as a pre-filter to exclude extreme outliers before shift detection, and 2) as a standalone safeguard for identifying early or localized system failures. For instance, [82] shows that anomaly detection improves the reliability of healthcare AI systems by filtering out unreliable inputs and preventing erroneous predictions.
@@ 6.1 Performance-Based Methods
These methods detect model drift by tracking changes in metrics such as accuracy, error rate, F1 score, or ROC-AUC using labeled data [52]. They compare performance across time windows or batches to identify significant degradation.
A classic example is [83], which monitors online error rates and triggers drift alarms when errors exceed a threshold, prompting model retraining or recalibration. This model-agnostic method was applied by Rotalinti [84] to clinical prediction models using the Wilcoxon test on ROC-AUC over time, showing that even high-performing models can experience silent drift in dynamic healthcare settings.
@@ 6.2 Model Output-Based Methods
These methods detect drift by tracking shifts in predicted outputs (e.g., softmax probabilities or confidence scores) without relying on ground-truth labels [52, 85, 86]. Such outputs act as compact proxies for model behavior, enabling scalable, label-free, and model-agnostic monitoring, which is particularly valuable in healthcare, where labels are often delayed or unavailable.
@@ 6.2.1 Distance-Based Output Monitoring
These methods are especially effective in multi-class settings, capturing subtle class-specific shifts through the full output vector and offering deeper insight into model behavior.
@@ 6.2.2 Score-Based Output Monitoring
The study [88] uses maximum softmax probability to flag prediction errors and distribution shifts. Evaluated on both vision and text datasets, the method shows that misclassified or OOD inputs tend to have lower softmax scores, making it a widely adopted baseline for OOD and drift detection. Energy-based methods [76] extend this idea by computing a log-sum-exp energy score from logits, which captures overall activation levels and is sensitive to subtle uncertainties.
@@ 6.3 Auxiliary Model-Based Method
These methods are modular and flexible, preserving the primary model’s integrity while enabling auxiliary modules to be trained/adapted independently to dynamic environments. The drawback is that extra computation is typically needed.
@@ 6.4 Calibration Drift Detection
Calibration drift occurs when a model’s predicted probabilities no longer align with actual outcome frequencies. For example, if a model predicts a class with 0.8 confidence, that prediction should be correct about 80% of the time. In dynamic clinical settings, this alignment can degrade over time.
@@ 7.2 Shift Type Disentanglement Method
Accurate identification of the specific type of distribution shift is essential for effective RCA. The study [94] proposes a framework that differentiates between label shift, covariate shift, and hybrid shift. The approach uses model output distributions to detect label shifts and leverages feature-based analysis via self-supervised encoders to detect covariate shifts. The combined “Duo detector” which integrates both techniques, achieves superior performance across various shift types. Evaluated on large-scale datasets from chest X-rays, mammography, and retinal fundus images, the method demonstrates strong generalizability.
The key takeaway is that each shift type necessitates tailored detection and correction strategies. For example, label shift often requires output reweighting, whereas covariate shift is better addressed through feature-level adjustments.
@@ 7.3 Causal Inference-Based Method
7.4 Model-Centric Analysis of Internal Dynamics
Another perspective on RCA involves examining the internal parameters of neural networks. For example, WeightWatcher [98] evaluates the spectral properties of weight matrices across layers in pretrained deep networks. Using norm-based and power-law-based metrics grounded in heavy-tailed self-regularization theory, WeightWatcher can detect when a model exhibits capacity loss, over-parameterization, or under-training. Such approaches highlight the value of incorporating layer-wise diagnostics into RCA, offering a model-centric complement to data-centric or causal inference methods.
7.5 Illustrative Example
Consider a clinical decision support model for pneumonia detection from chest X-rays. Over time, its accuracy drops when deployed in a new hospital. An RCA pipeline would proceed as follows:
  1. 1.
    Detection: Model monitoring identifies a significant performance drop for a specific patient subgroup (e.g., elderly patients).
  2. 2.
    Data-level analysis: Data monitoring reveals a shift in X-ray device type and contrast distribution compared to the training site.
  3. 3.
    Model-level analysis: Inspection of internal activations shows reduced feature sensitivity, i.e., a weakened correlation between intermediate layer features and the classification output, in mid-level convolutional layers.
  4. 4.
    Correction: Domain adaptation or calibration adjustment restores performance.
This example illustrates how RCA integrates data- and model-level analyses to identify the underlying cause of degradation and guide corrective action, which is essential for maintaining safe and reliable medical AI systems.
@@ 7.5 Illustrative Example
Table 1: Quantitative comparison of representative methods across the Detection-Diagnosis-Correction framework.
Stage Method Family (Examples) Prerequisites Cost
Detection Statistical distance measures (MMD, JS) Reference data, feature representations Medium-High
Statistical test (KL test, Chi-Square Test) Reference data, feature/output distributions Low
Supervised methods (Domain classifier) Access to cross-site data, model training Medium-High
Unsupervised methods (Autoencoder) Access to source data, model training Medium
Feature-based methods (SHAP) Access to model internals/predictions, labeled samples High
Out-of-Distribution Detection (MSP) Reference data, model outputs or feature embeddings Medium
Performance analysis (Accuracy) Labeled target data Low
Diagnosis LLM-based diagnosis Access to LLM resources, reference data High
Shift type disentanglement Reference data, labeled samples, model outputs Medium-High
Causal inference Labeled data with causal variables, model access High
Correction Unsupervised domain adaptation Source and target data, model access High
Test‐time adaptation Model access, unlabeled stream Medium-High
Re-training and Fine-Tuning Labeled data, model training High
Continual learning Sequential data, model update access Medium-High
@@ 8.1.1 Unsupervised Domain Adaptation (UDA)
It assumes access to labeled data in the source domain and unlabeled data in the target domain [100]. The goal is to align the latent representations across domains without requiring labels from the target side.
In the study [59], UDA is implemented to correct for data drift in COVID-19 detection models using audio data. After drift detection, UDA minimizes the MMD between the source (development) and target (post-development) data distributions. This process aligns the feature representations of the two domains. Experimental result validates UDA’s potential to maintain robust performance in evolving healthcare scenarios.